a new machine can make 10,000 aluminum cans three times faster than an older machine. with both machines working, 10,000 cans be made in 9h. how long would it take the new machine, working alone, to make the 10,000 cans?
Answer by mananth(16946)

older machine x cans /hour
new machine 3x cans /hour
..
both working ---10,000 cans - 9 hours
1 hour 10000/9 = 1111 cans /hour
3x+x= 1111 ( rounded off)
4x=1111
x= 1111/4
x= 277 cans /hour
new machine 3*277 = 831